# Demystify Protein Generation with Hierarchical Conditional Diffusion Models

### v1

@@ Abstract
Generating novel and functional protein sequences is critical to a wide range of applications in biology. Recent advancements in conditional diffusion models have shown impressive empirical performance in protein generation tasks. However, reliable generations of protein remain an open research question in de novo protein design, especially when it comes to conditional diffusion models. Considering the biological function of a protein is determined by multi-level structures, we propose a novel multi-level conditional diffusion model that integrates both sequence-based and structure-based information for efficient end-to-end protein design guided by specified functions. By generating representations at different levels simultaneously, our framework can effectively model the inherent hierarchical relations between different levels, resulting in an informative and discriminative representation of the generated protein. We also propose a Protein-MMD, a new reliable evaluation metric, to evaluate the quality of generated protein with conditional diffusion models. Our new metric is able to capture both distributional and functional similarities between real and generated protein sequences while ensuring conditional consistency. We experiment with the benchmark datasets, and the results on conditional protein generation tasks demonstrate the efficacy of the proposed generation framework and evaluation metric.
@@ 1 Introduction
Designing proteins with specific biological functions is a fundamental yet formidable challenge in biotechnology. It benefits wide-ranging applications from synthetic biology to drug discovery (Watson et al., 2023; Bose et al., 2016; Huang et al., 2016; Feng et al., 2024; Huang et al., 2024b, b; Lin et al., 2023). The challenge arises from the intricate interplay between protein sequence, structure, and function, which has not yet been fully understood (Dill et al., 2008). Traditional methods, such as directed evolution, rely on labor-intensive trial-and-error approaches involving random mutations and selective pressures, making the process time-consuming and costly (Arnold, 1998). Recently, generative models have emerged as promising tools for protein design, enabling the exploration of vast sequence-structure-function landscapes (Anand & Achim, 2022; Fu et al., 2024; Dauparas et al., 2022; Trippe et al., 2023). However, existing generative models—including those focused on enzyme engineering, antibody creation, and therapeutic protein development—are typically task-specific and require retraining for new design objectives (Fu et al., 2024; Dauparas et al., 2022). These limitations impede their adaptability and scalability across different protein families.
@@
While conditional generative models offer an end-to-end solution by directly linking the design process to the guidance, these models have been applied to protein generation (Komorowska et al., 2024; Klarner et al., 2024; Gruver et al., 2023). In conditional protein generation tasks, maintaining conditional consistency across diverse contexts and ensuring functional relevance are critical (Trippe et al., 2022; Hu et al., 2024). Specifically, the generated proteins should fully adhere to the specified functional constraints (Gretton et al., 2012). At the same time, achieving diversity and novelty in generated proteins is essential for successful design. In the literature, structural novelty can be assessed using Foldseek (van Kempen et al., 2022), which performs rapid protein structure searches against databases like PDB (Berman et al., 2000) and AlphaFold (Jumper et al., 2021) to ensure the generated proteins are novel compared to known structures. Diversity is measured using TM-score (Zhang & Skolnick, 2005), which calculates structural variation between the generated proteins themselves and between the generated and wild-type proteins (Hu et al., 2024).
@@
Despite the success of existing diffusion models in protein generation, these models only generate the protein representation at a single level and ignore hierarchical relations among different levels of representations. Choosing the level of granularity at which representing the comprehensive information of the protein raises significant concerns about the reliability of generated proteins in real-world applications. Motivated by the need to capture both the structural and functional nuances of protein sequences, we propose a novel multi-level conditional generative diffusion model for protein design that integrates both sequence-based (Lin et al., 2023) and structure-based (Wang et al., 2022a) hierarchical information. Specifically, our proposed method generates the protein at three different levels: the amino acid level, the backbone level, and the all-atom level. Generation at multi-levels enables efficient end-to-end generation of proteins with specified functions and modeling the inherent hierarchical relations between different representations, resulting in an informative and discriminative representation of the protein. Also, the conditional diffusion flow in the architecture preserves the hierarchical relations between different levels. Intuitively, a representation at the lower level (e.g., the atom level) can decide the potential representation space at the higher level (e.g., the amino acid level). Modeling such hierarchical relations can guarantee consistency at different levels. Our model incorporates a rigid-body 3D rotation-invariant preprocessing step combined with an autoregressive decoder to maintain SE(3)-invariance, ensuring accurate modeling of protein structures in 3D space. Figure 1 shows the proteins generated by different methods with the same input. The thin line indicates that the sequence is unlikely to undergo meaningful folding into a stable 3D structure. Compared with the baselines, our method can generate discriminative and functional proteins.
@@
We remark that it is still unknown how to assess the conditional consistency (Gretton et al., 2012) in de novo protein design. Specifically, the fundamental problem of properly evaluating conditional consistency is quantifying to what extent the generated protein adheres to the specified functional constraints. Unlike computer vision, where metrics such as FID (Heusel et al., 2017) have become a standard for assessing generated images, it is unclear whether such metrics are suitable for protein generation tasks. In protein design, the generated output cannot be as easily visualized or assessed as in images, making the choice of evaluation metrics even more critical. Therefore, how to adapt metrics like FID or Maximum Mean Discrepancy (Gretton et al., 2012) presents challenges. To address the challenges of evaluating the conditional consistency, we propose Protein-MMD, a metric based on Maximum Mean Discrepancy (MMD), to better capture both distributional and functional similarities between real and generated protein sequences, while ensuring conditional consistency. We prove that our Protein-MMD provides a more accurate measure that reflects the given condition. Experiments demonstrate that our proposed model outperforms existing approaches in generating diverse, novel, and functionally relevant proteins.
Our main contributions are summarized as follows:
@@
  • We design a novel multi-level conditional generative diffusion model that integrates sequence-based and structure-based information for efficient end-to-end protein design. Our framework incorporates a rigid-body 3D rotation-invariant preprocessing step to maintain SE(3)-invariance, modeling the hierarchical relations among different levels.
@@
  • We highlight the limitations of current evaluation metrics in protein generation tasks, particularly in conditional settings, and propose Protein-MMD, a novel metric to evaluate conditional consistency for protein generation tasks.
@@
2 Preliminary
2.1 Diffusion Models
Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020) are a class of generative models that learn to model data distributions. The forward diffusion process gradually adds noise to the data:
$q(x_{t}\mid x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I),$ (1)
where $x_{0}$ is the original data, $x_{t}$ is the data at timestep $t$, $\beta_{t}$ is the variance schedule, and $\mathcal{N}$ denotes a Gaussian distribution. The reverse diffusion process aims to model the posterior $q(x_{t-1}\mid x_{t})$, which is approximated by a neural network $\epsilon_{\theta}$ parameterized with $\theta$:
$p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t)),$ (2)
where $\mu_{\theta}$ and $\Sigma_{\theta}$ are functions of $x_{t}$ and timestep $t$. The model is trained to minimize the variational bound on negative log-likelihood:
$L_{\text{simple}}=\mathbb{E}_{x_{0},\epsilon,t}\left[\left|\epsilon-\epsilon_{\theta}(x_{t},t)\right|^{2}\right],$ (3)
where $t$ is uniformly sampled from ${1,\dots,T}$. Instead of using traditional U-Net architectures, recent works (Peebles & Xie, 2023a) leverage Transformer to predict the noise in the latent space (Rombach et al., 2022):
$\epsilon_{\theta}(z_{t},t)=\text{Transformer}(z_{t},t;\theta),$ (4)
where $z$ is a latent representation of $x$ and $\theta$ includes the parameters of the Transformer model.
2.2 Fréchet Inception Distance
Fréchet Inception Distance (FID) (Heusel et al., 2017) is commonly used to compute the Fréchet distance between feature representations of real and generated samples, extracted from a pre-trained network, such as Inception v3 (Szegedy et al., 2016). The FID between two distributions $p_{r}$ (real data) and $p_{g}$ (generated data) is computed by modeling the features as multivariate Gaussian distributions with means $\mu_{r},\mu_{g}$ and covariances $\Sigma_{r},\Sigma_{g}$:
$\text{FID}(p_{r},p_{g})=\|\mu_{r}-\mu_{g}\|^{2}_{2}+\text{Tr}\left(\Sigma_{r}+\Sigma_{g}-2(\Sigma_{r}\Sigma_{g})^{1/2}\right),$ (5)
where $\|\cdot\|_{2}$ denotes the Euclidean distance, and Tr is the trace operator. FID provides a measure of how close the real and generated data are, with lower FID values indicating high similarity.
Figure 2: The architecture of the multi-level diffusion model.
@@ 3.1 Multi-level Diffusion
Amino Acid Level Representation. As the 3D conformation dictates biochemical interactions (Huang et al., 2016; Dill et al., 2008), we first represent a protein’s structure as a graph $\mathcal{G}_{a}=(\mathcal{V}_{a},\mathcal{E}_{a})$, where $\mathcal{V}_{a}$ is the set of nodes corresponding to residues (amino acids), and $\mathcal{E}_{a}$ is the set of edges representing interactions between two residues. Specifically, an edge between two nodes $v_{i}$ and $v_{j}$ is established if the Euclidean distance between their $\text{C}\alpha$ atoms in 3D space is below a certain threshold, indicating a potential biochemical or structural interaction. At the amino acid level, each node $v_{i}\in\mathcal{V}_{a}$ corresponds to an amino acid and is represented by a vector $v_{i}=(\phi_{i};h_{i})$, where $\phi_{i}\in\mathbb{R}^{3}$ denotes the spatial coordinates of the amino acid’s $\text{C}\alpha$ atom in three-dimensional space, and $h_{i}$ abstracts biochemical or structural properties. Each edge is represented as an embedding of the sequential distance (Zhang et al., 2023).
@@
Backbone Level Representation. An amino acid consists of backbone atoms and side chain atoms. Similarly, we use backbone atom ($C$, $N$, $C_{\alpha}$) coordinates as the feature of in node of the backbone $\mathcal{V}_{b}$. We follow (Zhang et al., 2023) to compute three Euler angles $\tau_{i,j}^{1}$, $\tau_{i,j}^{2}$, $\tau_{i,j}^{3}$ between two backbone atoms $i$ and $j$. The angles will be integrated with the sequential distance as the edge feature. Backbone-level representation derives finer-grained protein information. With the three angles, the orientation between any two backbone planes can be determined to capture the backbone structures.
@@
Atom Level Representation. Atom-level representation considers all atoms in the protein and provides the most fine-grained information. There are several methods to treat an atom as a node in the representation (Hermosilla et al., 2021; Jing et al., 2021). Side chain torsion angles are important properties of protein structures (Jumper et al., 2021). In this paper, we also consider geometric representation at the atom level by incorporating the first four torsion angles: $\chi_{i}^{1}$, $\chi_{i}^{2}$, $\chi_{i}^{3}$, and $\chi_{i}^{4}$. With the complete geometric representation at the atom level, the diffusion model can capture 3D information about all atoms in a protein and distinguish any two distinct protein structures in nature.
@@
Encoding. We adopt a graph neural networks model (Wang et al., 2022b) to encode the representing at different levels by leveraging the message-passing mechanism. In many models dealing with the spatial positions of amino acids, SE(3)-equivariance is often leveraged to ensure the invariance of operations such as translation and rotation (Bose et al., 2016; Yim et al., 2023). We also introduce a novel method to ensure SE(3)-invariance by transforming each amino acid’s coordinates $\phi$. This step is crucial for facilitating the subsequent autoregressive decoding.
@@
where $\theta$ is the rotation angle between a node $v$ and the $x$-axis, and ${K}$ is the skew-symmetric matrix derived from the cross-product of $v$ and the unit vector along the $x$-axis. The third amino acid is rotated around the $x$-axis to place it in the positive $xy$-plane:
@@
The decoder at each level is an autoregressive Transformer (Vaswani, 2017) model that reconstructs the protein at each respective level. The autoregressive decoder can then use these transformed embeddings to reconstruct the information of a protein. At the sequence level, the decoder predicts the next amino acid token in the sequence. At the backbone level and the atom level, the decoder predicts geometric features (e.g., bond angles and distances) in an autoregressive fashion of each amino acid in the protein chain. Our method facilitates the use of SE(3)-invariant embeddings within an autoregressive framework. The decoder’s autoregressive nature allows it to progressively predict amino acid positions by leveraging the SE(3)-invariant representation.
@@
Hierarchical Diffusion with Conditional Flow: To achieve control over the conditional generation of proteins at multiple levels, we employ a novel hierarchical diffusion model with a conditional flow mechanism. This design enables fine-grained manipulation of protein structure generation under specific conditions, such as targeted functional attributes. The diffusion process is split into three distinct levels: all-atom, backbone, and amino acid (sequence). Conditional information is injected from a lower level to ensure conditional consistency throughout the generation process.
@@
$z^{i}_{t}=\epsilon^{i}(z^{i}_{t-1};z^{i-1}_{t}W^{i},c,\gamma_{t}),$ (22)
@@ 3.2 Evaluation of Conditional Consistency
$\vspace{-1mm}\text{MRR}=\frac{1}{|Q|}\sum_{i=1}^{|Q|}\frac{1}{\text{rank}_{i}},\\ \text{NMR}=\frac{1}{|Q|}\sum_{i=1}^{|Q|}\frac{\text{rank}_{i}-1}{N-1}.$ (24)
@@
In this paper, we propose Protein-MMD, a new evaluation metric that calculates the Maximum Mean Discrepancy (MMD) based on protein embeddings. Specifically, both real and generated protein sequences are encoded using the ESM2 language model (Lin et al., 2023), which provides biologically informed embeddings. ESM2 was chosen due to its ability to capture both structural and functional properties of proteins, thanks to its pretraining on a large protein corpus. This makes ESM2 particularly effective for evaluating distributional and functional similarities between real and generated proteins, aligning with the goals of de novo protein design:
[…]

### v2

# AI for Science

@@ Abstract
Generating novel and functional protein sequences is critical to a wide range of applications in biology. Recent advancements in conditional diffusion models have shown impressive empirical performance in protein generation tasks. However, reliable generation of proteins remains an open research question in de novo protein design, especially when it comes to conditional diffusion models. Considering the biological function of a protein is determined by multi-level structures, we propose a novel multi-level conditional diffusion model that integrates both sequence-based and structure-based information for efficient end-to-end protein design guided by specified functions. By generating representations at different levels simultaneously, our framework can effectively model the inherent hierarchical relations between different levels, resulting in an informative and discriminative representation of the generated protein. We also propose Protein-MMD (Maximum Mean Discrepancy), a new reliable evaluation metric, to evaluate the quality of generated protein with conditional diffusion models. Our new metric is able to capture both distributional and functional similarities between real and generated protein sequences while ensuring conditional consistency. Using conditional protein generation tasks with benchmark datasets, we demonstrate the efficacy of the proposed protein generation framework and evaluation metric.
Anonymous
@@ 1 Introduction
Designing proteins with specific biological functions is a fundamental yet formidable challenge in biotechnology. It benefits wide-ranging applications from synthetic biology to drug discovery [watson2023novo, DBLP:conf/iclr/BoseAHFRLNKB024, huang2016coming, DBLP:conf/iclr/FengLJML24, DBLP:conf/iclr/Huang0ZZZZCW0Y24, DBLP:conf/iclr/Huang0ZZZZCW0Y24, lin2023evolutionary]. The challenge arises from the intricate interplay between protein sequence, structure, and function, which has not yet been fully understood [dill2008protein]. Traditional methods, such as directed evolution, rely on labor-intensive trial-and-error approaches involving random mutations and selective pressures, making the process time-consuming and costly [arnold1998design]. Recently, generative models have emerged as promising tools for protein design, enabling the exploration of vast sequence-structure-function landscapes [anand2022protein, fu2024latent, dauparas2022robust, DBLP:conf/iclr/TrippeYTBBBJ23]. However, existing generative models—including those focused on enzyme engineering, antibody creation, and therapeutic protein development—are typically task-specific and require retraining for new design objectives [fu2024latent, dauparas2022robust]. These limitations impede their adaptability and scalability across different protein families.
While conditional generative models offer an end-to-end solution by directly linking the design process to the guidance, these models have been applied to protein generation [DBLP:conf/iclr/KomorowskaMDVLJ24, DBLP:conf/icml/KlarnerRMDT24, DBLP:conf/nips/GruverSFRHLRCW23]. In conditional protein generation tasks, maintaining conditional consistency across diverse contexts and ensuring functional relevance are critical [trippe2022diffusion, hu2024cpdiffusion-ss]. Specifically, the generated proteins should fully adhere to the specified functional constraints [gretton2012kernel]. At the same time, achieving diversity and novelty in generated proteins is essential for successful design. In the literature, structural novelty can be assessed using Foldseek [van2022foldseek], which performs rapid protein structure searches against databases like PDB [berman2000protein] and AlphaFold [jumper2021highly] to ensure the generated proteins are novel compared to known structures. Diversity is measured using TM-score [zhang2005tm], which calculates structural variation between the generated proteins themselves and between the generated and wild-type proteins [hu2024cpdiffusion-ss].
@@
Despite the success of existing diffusion models in protein generation, these models only generate the protein representation at a single level and ignore hierarchical relations among different levels of representations. Choosing the level of granularity at which representing the comprehensive information of the protein raises significant concerns about the reliability of generated proteins in real-world applications. Motivated by the need to capture both the structural and functional nuances of protein sequences, we propose a novel multi-level conditional generative diffusion model for protein design that integrates both sequence-based [lin2023evolutionary] and structure-based [wang2022learning] hierarchical information. Specifically, our proposed method generates the protein at three different levels: the amino acid level, the backbone level, and the all-atom level. Generation at multi-levels enables efficient end-to-end generation of proteins with specified functions and modeling the inherent hierarchical relations between different representations, resulting in an informative and discriminative representation of the protein. Also, the conditional diffusion flow in the architecture preserves the hierarchical relations between different levels. Intuitively, a representation at the lower level (e.g., the atom level) can decide the potential representation space at the higher level (e.g., the amino acid level). Modeling such hierarchical relations can guarantee consistency at different levels. Our model incorporates a rigid-body 3D rotation-invariant preprocessing step combined with an autoregressive decoder to maintain SE(3)-invariance, ensuring accurate modeling of protein structures in 3D space. Figure 1 shows the proteins generated by different methods with the same input. The thin line indicates that the sequence is unlikely to undergo meaningful folding into a stable 3D structure. Compared with the baselines, our method can generate discriminative and functional proteins.
@@
We remark that it is still unknown how to assess the conditional consistency [gretton2012kernel] in de novo protein design. Specifically, the fundamental problem of properly evaluating conditional consistency is quantifying to what extent the generated protein adheres to the specified functional constraints. Unlike computer vision, where metrics such as FID [heusel2017gans] have become a standard for assessing generated images, it is unclear whether such metrics are suitable for protein generation tasks. In protein design, the generated output cannot be as easily visualized or assessed as in images, making the choice of evaluation metrics even more critical. Therefore, how to adapt metrics like FID or Maximum Mean Discrepancy [gretton2012kernel] presents challenges. To address the challenges of evaluating the conditional consistency, we propose Protein-MMD, a metric based on Maximum Mean Discrepancy (MMD), to better capture both distributional and functional similarities between real and generated protein sequences, while ensuring conditional consistency. We prove that our Protein-MMD provides a more accurate measure that reflects the given condition. Experiments demonstrate that our proposed model outperforms existing approaches in generating diverse, novel, and functionally relevant proteins. Our main contributions are summarized as follows:
@@
  • We design a novel multi-level conditional generative diffusion model that integrates sequence-based and structure-based information for efficient end-to-end protein design.
@@
  • We highlight the limitations of current evaluation metrics in protein generation and propose Protein-MMD, a novel metric to evaluate conditional consistency for protein generation.
@@ 2.1 Multi-level Diffusion
Figure 2: The architecture of the multi-level diffusion model.
@@
Amino Acid Level Representation. As the 3D conformation dictates biochemical interactions [huang2016coming, dill2008protein], we first represent a protein’s structure as a graph $\mathcal{G}_{a}=(\mathcal{V}_{a},\mathcal{E}_{a})$, where $\mathcal{V}_{a}$ is the set of nodes corresponding to residues (amino acids), and $\mathcal{E}_{a}$ is the set of edges representing interactions between two residues. Specifically, an edge between two nodes $v_{i}$ and $v_{j}$ is established if the Euclidean distance between their $\text{C}\alpha$ atoms in 3D space is below a certain threshold, indicating a potential biochemical or structural interaction. At the amino acid level, each node $v_{i}\in\mathcal{V}_{a}$ corresponds to an amino acid and is represented by a vector $v_{i}=(\phi_{i};h_{i})$, where $\phi_{i}\in\mathbb{R}^{3}$ denotes the spatial coordinates of the amino acid’s $\text{C}\alpha$ atom in three-dimensional space, and $h_{i}$ abstracts biochemical or structural properties. Each edge is represented as an embedding of the sequential distance [DBLP:conf/iclr/ZhangXJCLD023].
@@
Backbone Level Representation. An amino acid consists of backbone atoms and side chain atoms. Similarly, we use backbone atom ($C$, NN, $C_{\alpha}$) coordinates as the feature of in node of the backbone $\mathcal{V}_{b}$. We follow [DBLP:conf/iclr/ZhangXJCLD023] to compute three Euler angles $\tau_{i,j}^{1}$, $\tau_{i,j}^{2}$, $\tau_{i,j}^{3}$ between two backbone atoms $i$ and $j$. The angles will be integrated with the sequential distance as the edge feature. Backbone-level representation derives finer-grained protein information. With the three angles, the orientation between any two backbone planes can be determined to capture the backbone structures.
@@
Atom Level Representation. Atom-level representation considers all atoms in the protein and provides the most fine-grained information. There are several methods to treat an atom as a node in the representation [DBLP:conf/iclr/HermosillaSLFVK21, DBLP:journals/corr/abs-2106-03843]. Side chain torsion angles are important properties of protein structures [jumper2021highly]. In this paper, we also consider geometric representation at the atom level by incorporating the first four torsion angles: $\chi_{i}^{1}$, $\chi_{i}^{2}$, $\chi_{i}^{3}$, and $\chi_{i}^{4}$. With the complete geometric representation at the atom level, the diffusion model can capture 3D information about all atoms in a protein and distinguish any two distinct protein structures in nature.
@@
Encoding. We adopt a graph neural networks model [DBLP:conf/nips/WangLLLJ22] to encode the representing at different levels by leveraging the message-passing mechanism. In many models dealing with the spatial positions of amino acids, SE(3)-equivariance is often leveraged to ensure the invariance of operations such as translation and rotation [DBLP:conf/iclr/BoseAHFRLNKB024, DBLP:conf/icml/YimTBMDBJ23]. We also introduce a novel method to ensure SE(3)-invariance by transforming each amino acid’s coordinates $\phi$. This step is crucial for facilitating the subsequent autoregressive decoding.
@@
where $\theta$ is the rotation angle between a node $v$ and the $x$-axis, and K{K} is the skew-symmetric matrix derived from the cross-product of $v$ and the unit vector along the $x$-axis. The third amino acid is rotated around the $x$-axis to place it in the positive $xy$-plane:
@@
The decoder at each level is an autoregressive Transformer [vaswani2017attention] model that reconstructs the protein at each respective level. The autoregressive decoder can then use these transformed embeddings to reconstruct the information of a protein. At the sequence level, the decoder predicts the next amino acid token in the sequence. At the backbone level and the atom level, the decoder predicts geometric features (e.g., bond angles and distances) in an autoregressive fashion of each amino acid in the protein chain. Our method facilitates the use of SE(3)-invariant embeddings within an autoregressive framework. The decoder’s autoregressive nature allows it to progressively predict amino acid positions by leveraging the SE(3)-invariant representation.
@@
Hierarchical Diffusion with Conditional Flow: To achieve control over the conditional generation of proteins at multiple levels, we employ a novel hierarchical diffusion model with a conditional flow mechanism. This design enables fine-grained manipulation of protein structure generation under specific conditions, such as targeted functional attributes. The diffusion process is split into three distinct levels: all-atom, backbone, and amino acid (sequence). Conditional information is injected from a lower level to ensure conditional consistency.
@@ 2.2 Evaluation of Conditional Consistency
$\vskip-2.84526pt\text{MRR}=\frac{1}{|Q|}\sum_{i=1}^{|Q|}\frac{1}{\text{rank}_{i}},\\ \text{NMR}=\frac{1}{|Q|}\sum_{i=1}^{|Q|}\frac{\text{rank}_{i}-1}{N-1}.$ (19)
@@
In this paper, we propose Protein-MMD, a new evaluation metric that calculates the Maximum Mean Discrepancy (MMD) based on protein embeddings. Specifically, both real and generated protein sequences are encoded using the ESM2 language model [lin2023evolutionary], which provides biologically informed embeddings. ESM2 was chosen due to its ability to capture both structural and functional properties of proteins, thanks to its pretraining on a large protein corpus. This makes ESM2 particularly effective for evaluating distributional and functional similarities between real and generated proteins, aligning with the goals of de novo protein design:
$\text{Protein-MMD}(p_{r},p_{g})=\left\|\frac{1}{n}\sum_{i=1}^{n}\varphi(x_{i})-\frac{1}{m}\sum_{j=1}^{m}\varphi(y_{j})\right\|_{\mathcal{H}}^{2},$ (20)
@@
To validate the effectiveness of Protein-MMD and other metrics, we apply the evaluation metrics on the Enzyme Commission (EC) dataset, which categorizes proteins based on the reactions they catalyze using EC numbers. We focus on seven classes from the first EC number category for our conditional generation task. In Table 1, we compare three evaluation metrics: MMD (considering only sequence statistics as presented in [kucera2022conditional]), Protein-FID (using ESM2 in place of Inception for protein generation), and Protein-MMD. All metrics are used to compute the Accuracy, Mean Reciprocal Rank (MRR), and Normalized Mean Rank (NMR) scores to evaluate how it performs in evaluating the conditional consistency. As observed, Protein-MMD outperforms both MMD and Protein-FID across all evaluation metrics. The higher accuracy and MRR scores indicate that Protein-MMD better captures the conditional consistency of the proteins in the test set. The lower NMR score further demonstrates that Protein-MMD ranks the correct target class higher in comparison to other metrics, validating its effectiveness in conditional protein generation tasks. While Protein-MMD proves to be the most effective metric according to our framework, we acknowledge the widespread use of FID in generative modeling tasks. Therefore, we will continue to report Protein-FID results alongside Protein-MMD in subsequent experiments.
3 Experiments
3.1 Experimental Setup
We compared our model against several baselines, each representing distinct approaches to protein generation. ProteoGAN [kucera2022conditional] is a GAN-based method, while ESM2 [lin2023evolutionary] and ProstT5 [heinzinger2023prostt5] are Transformer-based language models specifically designed for protein sequence modeling. ProteinMPNN [dauparas2022robust] and LatentDiff [fu2024latent], on the other hand, are graph-based models, with LatentDiff also incorporating a diffusion-based framework, specifically using a latent diffusion approach. For each model, we evaluate the performance using both diversity metrics (TM-score, RMSD, and Seq.ID) and conditional consistency metrics (Protein-MMD and Protein-FID). Higher RMSD, lower TM-score, and lower Seq.ID indicate higher diversity, while lower Protein-MMD and Protein-FID values signify higher conditional consistency between the generated and real protein distributions. More detailed settings can be found in Appendix A.1.